Using the formula for the length of the arc, compute s, r or q which ever is missing.

s = 10pi r = 8


1
Expert's answer
2021-01-26T03:45:11-0500

"2\pi r=8."

"r=\frac{8}{2\pi}=\frac{4}{\pi}."